OpenStudy (anonymous):

What is the increasing interval of x^2-6x=5

OpenStudy (kainui):

Did you type the equation correctly?

OpenStudy (anonymous):

Yes \[x ^{2}-6x+5\]

OpenStudy (anonymous):

can you tell me what the increasing interval of x^2 is?

OpenStudy (anonymous):

is it (0,infinity)

OpenStudy (anonymous):

\[\left( 0,\infty \right)\]

OpenStudy (anonymous):

that's right now you must complete the square on what you have: f(x) = x^2 -6x - 5 this is equivalent to: \[(x-3)^2 - 14 \] \[(x-h)^2 + k\]is the same graph as x^2 except it is shifted h units to the right and k units up so the graph is: